OpenStudy (anonymous):

An open box constructed from a tin sheet has a square base and a volume of 216 in^3. Find the dimensions of the box, assumng that the minimum amount f material was used in its construction

OpenStudy (anonymous):

i have no clue where to go

OpenStudy (anonymous):

this is optimization btw

OpenStudy (anonymous):

put x = length of square base. then the base has area \[x^2\]

OpenStudy (anonymous):

the height is h but the volume is \[x^2\times h = 216\] solving for h gives \[h=\frac{216}{h}\]

OpenStudy (anonymous):

now the surface area is \[x^2 + 4xh \] yes?

OpenStudy (anonymous):

area of base plus 4 times area of sides

OpenStudy (anonymous):

oh damn i made a typo. solving for h gives \[h=\frac{216}{x^2}\]

OpenStudy (anonymous):

no the nonsense i wrote.

OpenStudy (anonymous):

sorry i slipped away from the computer

OpenStudy (anonymous):

so all in all we get a surface area of \[A(x)=x^2+4\times x \times \frac{216}{x^2}\] \[A(x)=x^2+\frac{54}{x}\]

OpenStudy (anonymous):

thats ok i made a stupid typo

OpenStudy (anonymous):

surface area is same as volume right...

OpenStudy (anonymous):

oh heck no

OpenStudy (anonymous):

volume is length times width times hight. 3 dimensions

OpenStudy (anonymous):

sorry can you hold on for a sec while i try to catch up to what your doing please

OpenStudy (anonymous):

surface area is an area. 2 dimensions. i wait. ignore the line where i wrote \[h=\frac{216}{h}\] it should be \[h=\frac{216}{x^2}\]

OpenStudy (anonymous):

oh ok lol sorry i just saw it the sites really laggy

OpenStudy (anonymous):

true that

OpenStudy (anonymous):

how'd u get the surface area?

OpenStudy (anonymous):

main goal is obviously to find a function that gives the surface area. then we minimize that function

OpenStudy (anonymous):

ok surface area

OpenStudy (anonymous):

base is a square with side x, so area is \[x^2\]

OpenStudy (anonymous):

4 sides have base x and height h so surface area of each side is \[xh\]

OpenStudy (anonymous):

there are 4 of them so we have all together \[x^2+4xh\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

area of base plus 4 times area of sides. but we also have a "constraint' namely that the volume is 216 cubic whatevers

OpenStudy (anonymous):

inches lol

OpenStudy (anonymous):

so we can solve for h in terms of x via \[V= 216 = x^2\times h \]

OpenStudy (anonymous):

making \[h=\frac{216}{x^2}\]

OpenStudy (anonymous):

oh ok and thats how u got the x^2 + 4*x*216/x^2

OpenStudy (anonymous):

now we plug that back in to the formula for the total surface area to get a function of one variable namely \[A(x)=x^2+\frac{54}{x}\]

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

:) lol it just hit me... sorry

OpenStudy (anonymous):

oh wait you have it and i made a mistake

OpenStudy (anonymous):

yeah the numerator right? of the 2nd term

OpenStudy (anonymous):

\[\frac{4\times 216\times x }{x^2}=\frac{864}{x}\]

OpenStudy (anonymous):

i divided instead of multiplied, brain fart

OpenStudy (anonymous):

so now we have to find the derivative

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so finally, assuming i did not make another bone head mistake, we have \[A(x)=x^2+\frac{864}{x}\]

OpenStudy (anonymous):

2x-864x^-2

OpenStudy (anonymous):

derivative is not to much work here. \[\frac{d}{dx} A(x)=x^2-\frac{864}{x^2}\]

OpenStudy (anonymous):

2x yes

OpenStudy (anonymous):

\[\frac{d}{dx}A(x)=2x -\frac{864}{x^2}\]

OpenStudy (anonymous):

set this beast = 0 and solve

OpenStudy (anonymous):

i got the same thing i just could hav took my x back to the bottom

OpenStudy (anonymous):

almost do it in your head. numerator is \[2x^3-864\]

OpenStudy (anonymous):

there is a critical point at 0 but we can ignore it because it makes no sense to have a box with base 0

OpenStudy (anonymous):

yay! im doing the same thing

OpenStudy (anonymous):

\[2x^3-864=0\] \[x^3=432\] \[x=\sqrt[3]{463}\]

OpenStudy (anonymous):

whatever that is

OpenStudy (anonymous):

7.736 rounded

OpenStudy (anonymous):

how's we do?

OpenStudy (anonymous):

7.56 its 434

OpenStudy (anonymous):

yeah well it is late

OpenStudy (anonymous):

half of 864 =432 so we are both wrong!

OpenStudy (anonymous):

7356 rounded

OpenStudy (anonymous):

7.56 i mean

OpenStudy (anonymous):

hmm yes i have it on my scraps and im typing the wrong thing

OpenStudy (anonymous):

you got the right answer in any case. good work!

OpenStudy (anonymous):

thanks you too as always but do i find my minimum by plugging in x=0, x = 7.56 what are my endpoints?